The positive difference between x and 2/9 is the same as the positive difference between 3/7 and 4/11. what could be the value of x?
Answer by edjones(8007)

3/7 - 4/11
33/77 - 28/77 = 5/77
x-(2/9)=5/77
x=(5/77)+(2/9)
=45/693 + 154/693
= 199/693
.
Ed